A 6% coupon bearing bond pays interest semi-annually and has a maturity of 21 years. If the current price of the bond is $1,074.49, what is the yield to maturity of this bond? (Answer to the nearest hundredth of a percent, e.g. 12.34%)
Answer #1

Number of periods = 21 * 2 = 42

Coupon = (0.06 * 1000) / 2 = 30

Yield to maturity = 5.40%

Keys to use in a financial calculator: 2nd I/Y 2, FV 1000, PV -1,074.49, N 42, PMT 30, CPT I/Y